ΔE Δt ≃ h̄

ΔEΔt ≃ h̄

\frac{Δλ}{λ_{0}} = \sqrt{\frac{1 − (v/c)}{1 + (v/c)}}

\frac{Δλ}{λ_{0}} ≃ \frac{v}{c}

mean free path ℓ = \frac{1}{nσ}

v_{thermal} = \sqrt{\frac{2kT}{m}}

Δt = \frac{ℓ}{v_{thermal}} = \frac{1}{nσ} \sqrt{\frac{m}{2kT}}